find the inverse and check.
h(x)= 1/3 x + 1
h(x)= 7/8x-2
d(x)= 5x+7/2x+4

Answer by jankiz(13)

Always remember, the inverse is a way to undo the equation.
First problem: H(x)=1/3x+1. Which is y=1/3x+1. You subtract 1 from both sides, then you multiply both sides by three. So the inverse is: 3(y-1).
H(x)=7/8x-2. Which is y=7/8x-2. First yo add 2 to both sides, then multiply by the reciprocal of 8/7. You get 8/7(y+2).
D(x)=5x+7/2x+4. Combine like terms. So, y=8.5x+4. Subtract 4 from both sides, then divide by 8.5. So, (y-4)/(8.5).